OpenStudy (anonymous):

Marna wants to store 200 cubic inches of homemade pasta sauce in a food storage container. Should she choose the cylinder with a radius of 4 inches and a height of 6 inches, or the cube with a side length of 5 inches? Explain.

OpenStudy (anonymous):

cube volume=side^3=5^3=125 not enough.

OpenStudy (anonymous):

cylinder volume=2*pi*r*h=2*4*6*pi=48*pi=3,14*48=150,something

OpenStudy (anonymous):

if you made no typos, the use of the cube is preferred, since the container fills at exactly 150ci

OpenStudy (anonymous):

125*

OpenStudy (anonymous):

so the cube

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

the cylinder is too much